# Structure-Informed Estimation for Pilot-Limited MIMO Channels via Tensor Decomposition

### v1

@@ Abstract
Channel estimation in wideband multiple-input multiple-output (MIMO) systems faces fundamental pilot overhead limitations in high-dimensional beyond-5G and sixth-generation (6G) scenarios. This paper presents a hybrid tensor–neural architecture that formulates pilot-limited channel estimation as low-rank tensor completion from sparse observations—a fundamentally different setting from prior tensor methods that assume fully observed received signal tensors. A canonical polyadic (CP) baseline implemented via a projection-based scheme (Tucker completion under partial observations) and Tucker decompositions are compared under varying signal-to-noise ratio (SNR) and scattering conditions: CP performs well for specular channels matching the multipath model, while Tucker provides greater robustness under model mismatch. A lightweight three-dimensional (3D) U-Net learns residual components beyond the low-rank structure, bridging algebraic models and realistic propagation effects. Empirical recovery threshold analysis shows that sample complexity scales approximately with intrinsic model dimensionality $L(N_{r}+N_{t}+N_{f})$ rather than ambient tensor size $N_{r}N_{t}N_{f}$, where $L$ denotes the number of dominant propagation paths. Experiments on synthetic channels demonstrate 10–20 dB normalized mean-square error (NMSE) improvement over least-squares (LS) and orthogonal matching pursuit (OMP) baselines at 5–10% pilot density, while evaluations on DeepMIMO ray-tracing channels show 24–44% additional NMSE reduction over pure tensor-based methods.
@@ 1 Introduction
Wideband multiple-input multiple-output (MIMO) systems with large antenna arrays are fundamental enablers of high-capacity wireless links in fifth-generation (5G) and beyond networks. Point-to-point MIMO configurations are particularly relevant for wireless backhaul, fixed wireless access, vehicle-to-infrastructure communications, and satellite feeder links [1, 2]. However, the spectral efficiency gains from spatial multiplexing critically depend on accurate channel state information (CSI), whose acquisition becomes progressively challenging as antenna counts, bandwidth, and mobility increase [3].
Recent advances in radio-frequency (RF) integration have fundamentally altered the architectural landscape of large-array systems. Fully digital transceivers—once considered impractical at millimeter-wave frequencies—are now recognized as scalable and energy-efficient, with commercial deployments already underway at sub-6 GHz and expected at millimeter-wave (mmWave) frequencies [4]. This architectural shift is further motivated by emerging sixth-generation (6G) use cases such as Integrated Sensing and Communications (ISAC), where fine-grained, per-antenna access to received signals is essential for joint optimization across communication and sensing tasks [5]. The present work focuses on fully digital MIMO orthogonal frequency-division multiplexing (MIMO-OFDM) architectures with direct baseband access to individual antenna-domain channel coefficients [^1] Hybrid beamforming architectures compress spatial information prior to baseband processing, limiting the flexibility required for ISAC. Fully digital transceivers preserve the complete spatial–frequency channel structure, naturally motivating estimation frameworks operating on the full antenna-domain tensor..
Conventional estimation techniques such as least squares (LS) and linear minimum mean-square error (LMMSE) require pilot resources scaling with channel dimensionality. In wideband MIMO-OFDM systems, the channel comprises $N_{r}\times N_{t}\times N_{f}$ complex coefficients, translating to tens of thousands of unknowns for systems with tens of antennas and hundreds of subcarriers [6, 7]. Deep learning addresses pilot scarcity by learning nonlinear mappings from sparse observations to full reconstructions [8, 9, 10, 11, 12], achieving strong performance when training and deployment conditions match. However, these methods typically require large labeled datasets, offer limited interpretability, and generalize poorly under distribution shifts—limitations particularly pronounced in pilot-limited regimes. A fundamental challenge is the limited availability of representative training data: unlike computer vision or natural language processing (NLP), acquiring labeled physical-layer data from operational networks is costly and often impractical [4].
@@
It is instructive to contrast tensor completion with compressed sensing (CS). CS-based approaches assume sparsity in a predefined angular–delay dictionary and recover coefficients from linear measurements, but are susceptible to basis mismatch in off-grid scenarios. Tensor completion instead exploits low-dimensional multilinear structure—such as low canonical polyadic (CP) or Tucker rank—directly in the native antenna–frequency domain without explicit dictionary construction. These perspectives are complementary: sparsity captures parametric specular structure, while multilinear low-rank models capture broader correlations and offer robustness under richer scattering conditions.
Prior tensor-based channel estimation has demonstrated substantial gains. Zhou et al. [17] formulated mmWave estimation as CP decomposition of a fully observed received signal tensor after RF combining—an overdetermined problem targeting hybrid architectures. Tensor methods have also been applied to MIMO relay systems via nested Tucker decompositions [19] and to time-varying channels under high mobility [20]. However, these formulations assume fully observed tensors and do not address pilot-limited completion.
The present work addresses a complementary scenario: pilot-limited estimation in fully digital systems where only a sparse subset of channel tensor entries is observed. This severely underdetermined inverse problem requires regularization through low-rank tensor completion—a fundamentally different mathematical setting applicable to a broader class of MIMO-OFDM systems [4].
The main contributions are:
@@
    Hybrid tensor–neural architecture: A two-stage estimator combining low-rank tensor completion with neural residual learning, achieving 24–44% NMSE reduction over pure tensor methods on DeepMIMO ray-tracing channels.
@@
    Tensor completion formulation for pilot-limited MIMO: Unlike prior methods assuming fully observed tensors [17], the problem is formulated as tensor completion from sparse pilot samples.
@@
    Comparative CP versus Tucker analysis: CP outperforms Tucker at moderate-to-high SNR for specular channels, while Tucker provides greater robustness under model mismatch.
  4. 4.
    Comprehensive validation: Evaluation against LS and OMP baselines demonstrates 10–20 dB NMSE improvement at 5–10% pilot density on both synthetic and ray-tracing channels.
@@ 2 System Model
Throughout this paper, scalars are denoted by lowercase letters (e.g., $\alpha$), vectors by bold lowercase letters (e.g., $\mathbf{a}$), matrices by bold uppercase letters (e.g., $\mathbf{A}$), and tensors by calligraphic letters (e.g., $\mathcal{H}$). The wideband MIMO channel is represented as a third-order tensor $\mathcal{H}\in\mathbb{C}^{N_{r}\times N_{t}\times N_{f}}$, where $N_{r}$, $N_{t}$, and $N_{f}$ denote the number of receive antennas, transmit antennas, and subcarriers, respectively. The $k$-th entry of a vector $\mathbf{w}$ is denoted by $\mathbf{w}[k]$. The outer product is denoted by $\circ$, the $n$-mode product by $\times_{n}$, the Frobenius norm by $\|\cdot\|_{F}$, the Hadamard (element-wise) product by $\odot$, and the Kronecker product by $\otimes$. All tensor dimensions and ranks are defined consistently with this convention throughout the paper.
A wideband MIMO–OFDM system is considered, with $N_{t}$ transmit antennas, $N_{r}$ receive antennas, and $N_{f}$ subcarriers. Let $\mathbf{H}[k]\in\mathbb{C}^{N_{r}\times N_{t}}$ denote the frequency-domain channel matrix at subcarrier $k\in\{1,\ldots,N_{f}\}$. The received signal is
@@
In practical pilot-aided OFDM training, the channel tensor is not directly observed. Pilot symbols yield local linear estimates at the antenna–subcarrier resource elements where pilots are transmitted. Accordingly, we model $\mathcal{P}_{\Omega}(\mathcal{H})$ as the collection of these pilot-based observations, i.e., noisy samples of the channel entries at the pilot locations. This abstraction reduces the training process to entry-wise sampling of the channel tensor (with noise), which is standard in tensor completion and isolates the impact of pilot sparsity from waveform-specific effects.
@@ 2.1 Pilot-Based Channel Estimation
During the training phase, known pilot symbols are transmitted over a subset of resource elements $(i,j,k)$ of the channel tensor, so the channel is observed only at a subset of its coefficients. Stacking the frequency-domain channel matrices $\mathbf{H}[k]\in\mathbb{C}^{N_{r}\times N_{t}}$ across all $N_{f}$ subcarriers yields the wideband MIMO channel represented as a third-order tensor
$\mathcal{H}\in\mathbb{C}^{N_{r}\times N_{t}\times N_{f}},$ (2)
whose three modes correspond to receive antennas, transmit antennas, and frequency, respectively.
Pilot-based acquisition provides access only to a subset of tensor entries indexed by
@@
The resulting observation model is [17]
@@
and $\mathcal{N}$ is i.i.d. complex Gaussian noise. The SNR is defined with respect to the average power of the channel entries at pilot locations (i.e., over $\Omega$), ensuring comparability across different pilot densities. Equivalently, $\mathcal{P}_{\Omega}(\mathcal{H})=\mathcal{M}\odot\mathcal{H}$, where $\mathcal{M}\in\{0,1\}^{N_{r}\times N_{t}\times N_{f}}$ is a binary mask with $\mathcal{M}]_{ijk}=1$ if $(i,j,k)\in\Omega$. The noise tensor $\mathcal{N}$ is obtained by stacking the per-subcarrier noise vectors $\mathbf{n}[k]$ across receive antennas and frequency indices, consistently with (1).
The tensor completion formulation models pilot-limited observation as entry-wise sampling over an index set $\Omega$. In the experiments, $\Omega$ is generated by a pilot mask with either random sampling, a regular grid, or a comb pattern in frequency. For theoretical discussion this work adopts the standard uniform random sampling model (corresponding to the random mask), while structured masks are treated as deterministic sampling patterns under the same observation model.
@@
only a small fraction of channel tensor coefficients is directly observed. In this pilot-limited regime, recovering the full channel tensor $\mathcal{H}$ from partial and noisy observations $\mathcal{Y}_{\Omega}$ constitutes a severely ill-posed inverse problem: unknowns vastly exceed measurements, the observation operator is highly rank-deficient, and the solution is non-unique and noise-sensitive [6, 7, 22]. This fundamental limitation motivates incorporating structural priors that exploit intrinsic spatial and spectral correlations of the channel to regularize the estimation problem and enable reliable reconstruction.
@@ 2.2 Multipath Channel Model
where $L$ denotes the number of dominant propagation paths, $\alpha_{\ell}\in\mathbb{C}$ are complex path gains, $\mathbf{a}_{r}(\theta_{\ell})$ and $\mathbf{a}_{t}(\phi_{\ell})$ are receive and transmit array steering vectors associated with angles of arrival and departure, respectively, and $\tau_{\ell}\in[0,1)$ represents the normalized propagation delay. The exponential term models frequency-dependent phase rotation induced by path delay across OFDM subcarriers.
Eq. (7) reveals that the wideband channel tensor $\mathcal{H}$ can be expressed as a superposition of rank-one components across spatial and frequency dimensions. Consequently, $\mathcal{H}$ admits an intrinsic CP tensor representation [17, 23], with effective tensor rank governed by the number of dominant propagation paths rather than ambient dimensionality. This physically grounded low-rank structure directly motivates tensor-based channel modeling and underpins the structure-informed estimation framework developed in the sequel.
@@ 3 Tensor Representations and Algebraic Priors
This section formalizes how fundamental propagation mechanisms in wideband MIMO channels manifest as intrinsic algebraic structure. Building on the multipath model introduced in Section 2, the channel is represented as a low-dimensional tensor whose multilinear structure reflects sparse specular propagation, array-induced spatial correlations, and frequency coherence due to finite delay spread. CP and Tucker decompositions are introduced as principled parameterizations of this structure, providing physically interpretable and computationally efficient priors for pilot-limited channel estimation. Importantly, the two decompositions embody different physical assumptions and offer complementary trade-offs between model compactness and representational flexibility.
3.1 Channel Tensor Representation
Rather than estimating each frequency-domain channel matrix $\mathbf{H}[k]$ independently, the wideband MIMO channel is modeled as a third-order tensor $\mathcal{H}\in\mathbb{C}^{N_{r}\times N_{t}\times N_{f}}$, whose modes correspond to receive antennas, transmit antennas, and subcarriers, respectively. This tensorial representation explicitly exposes correlations across spatial and spectral dimensions induced by array geometry and propagation physics, which per-subcarrier estimation approaches do not exploit [16, 17]. In time-varying scenarios, an additional temporal mode may be incorporated, yielding a higher-order tensor without altering the underlying algebraic structure.
@@ 3.2 Canonical Polyadic Decomposition
where $k$ indexes the subcarrier frequencies. Absorbing the complex gain $\alpha_{\ell}$ into the receive-mode factor $\mathbf{u}_{\ell}$ is a matter of convention; the gain could equivalently be absorbed into $\mathbf{v}_{\ell}$, into $\mathbf{w}_{\ell}$, or distributed across multiple factors. This scaling ambiguity is intrinsic to the CP decomposition: under the Kruskal uniqueness condition, the CP model is unique only up to permutation and scaling of its rank-one components [23].
This correspondence makes explicit that each rank-one tensor component represents a single physical propagation path, implying that the effective CP rank is governed by the number of dominant paths rather than the ambient channel dimensionality. As a result, CP is particularly well matched to propagation regimes dominated by specular ray-like components, such as millimeter-wave channels around 28 GHz, where a small number of strong paths typically capture most of the channel energy. This property underlies the effectiveness of CP-based models for sparse-channel and large-scale MIMO estimation [17]. It is worth emphasizing that the Kruskal uniqueness condition guarantees factor identifiability up to permutation and scaling, which is sufficient for reliable channel reconstruction. Extracting explicit physical parameters, such as angles of arrival or departure, from the estimated factors would require additional structural constraints—e.g., array calibration or explicit enforcement of Vandermonde structure—beyond the scope of this work, which focuses on CSI reconstruction rather than direct parameter estimation.
In the experiment, Tucker reconstruction is performed via alternating projections, consisting of truncated higher-order singular value decomposition (HOSVD) followed by enforcement of consistency with the observed entries. The baseline labeled “CP” corresponds, by default, to a projection-based variant that uses a stable Tucker-completion step under partial observations and returns the completed tensor; direct masked CP-Alternating Least Squares (ALS) is not used here due to numerical instability in sparse, noisy settings.

### v2

# arXiv:2602.04083v2[eess.SP] 02 Mar 2026

@@ Abstract
Accurate channel state information in wideband multiple-input multiple-output (MIMO) systems is fundamentally constrained by pilot overhead, a challenge that intensifies as antenna counts and bandwidths scale toward 6G. This paper proposes a structure-informed hybrid estimator that formulates pilot-limited MIMO channel estimation as low-rank tensor completion from sparse pilot observations—a severely underdetermined inverse problem that prior tensor approaches avoid by assuming fully observed received signal tensors. Canonical polyadic (CP) and Tucker decompositions are comparatively analyzed: CP excels for specular channels whose rank-one multipath structure matches the CP parameterization exactly, while Tucker provides greater numerical stability at extreme pilot scarcity where CP exhibits heavy-tail divergence. A lightweight 3D U-Net learns residual components beyond the dominant low-rank structure, compensating for diffuse scattering and hardware non-idealities that algebraic priors alone cannot capture. On synthetic specular channels, Tucker completion achieves $10.88$ dB NMSE improvement over least squares and $7.83$ dB over orthogonal matching pursuit at $\rho=10\%$ pilot density; CP outperforms Tucker by $13.11$ dB at SNR = 20 dB under the specular multipath model. On DeepMIMO ray-tracing channels, the hybrid estimator surpasses CP by $2.26$ dB and Tucker by $4.80$ dB at $\rho=8\%$, while remaining stable at $\rho=2\%$ where CP diverges; algebraic structure consistently outperforms unconstrained deep learning across the full pilot-density range, with a margin growing from $1.53$ dB at $\rho=2\%$ to $5.67$ dB at $\rho=20\%$. Empirical recovery threshold analysis confirms that sample complexity scales with intrinsic channel dimensionality—governed by the number of dominant propagation paths—rather than with the ambient tensor size.
@@ 1 Introduction
Wideband multiple-input multiple-output (MIMO) systems with large antenna arrays are fundamental enablers of high-capacity wireless links in 5G and beyond networks. Point-to-point MIMO configurations are particularly relevant for wireless backhaul, fixed wireless access, vehicle-to-infrastructure communications, and satellite feeder links [1, 2]. However, the spectral efficiency gains from spatial multiplexing critically depend on accurate channel state information (CSI), whose acquisition becomes progressively challenging as antenna counts, bandwidth, and mobility increase [3].
Recent advances in radio-frequency (RF) integration have fundamentally altered the architectural landscape of large-array systems. Fully digital transceivers—once considered impractical at millimeter-wave frequencies—are now recognized as scalable and energy-efficient, with commercial deployments already underway at sub-6 GHz and expected at millimeter-wave (mmWave) frequencies [4]. This architectural shift is further motivated by emerging 6G use cases such as Integrated Sensing and Communications (ISAC), where fine-grained, per-antenna access to received signals is essential for joint optimization across communication and sensing tasks [5]. The present article focuses on fully digital MIMO orthogonal frequency-division multiplexing (MIMO-OFDM) architectures with direct baseband access to individual antenna-domain channel coefficients.
Conventional estimation techniques such as least squares (LS) and linear minimum mean-square error (LMMSE) require pilot resources scaling with channel dimensionality. In wideband MIMO-OFDM systems, the channel comprises $N_{r}N_{t}N_{f}$ complex coefficients, where $N_{r}$, $N_{t}$, and $N_{f}$ denote the number of receive antennas, transmit antennas, and subcarriers, respectively, translating to tens of thousands of unknowns for systems with tens of antennas and hundreds of subcarriers [6, 7]. Deep learning addresses pilot scarcity by learning nonlinear mappings from sparse observations to full reconstructions [8, 9, 10, 11, 12], achieving strong performance when training and deployment conditions match. However, these methods typically require large labeled datasets, offer limited interpretability, and generalize poorly under distribution shifts—limitations particularly pronounced in pilot-limited regimes. A fundamental challenge lies in the limited availability of representative training data: unlike computer vision or natural language processing (NLP), acquiring labeled physical-layer data from operational networks is costly and often impractical [4].
@@
A comparison between tensor completion and compressed sensing (CS) highlights key structural differences. CS-based approaches assume sparsity in a predefined angular–delay dictionary and recover coefficients from linear measurements, but are susceptible to basis mismatch in off-grid scenarios. Tensor completion instead exploits low-dimensional multilinear structure—such as canonical polyadic (CP) or Tucker rank—directly in the native antenna–frequency domain without explicit dictionary construction. These perspectives are complementary: sparsity captures parametric specular structure, while multilinear low-rank models capture broader correlations and offer robustness under richer scattering conditions.
Prior tensor-based channel estimation has demonstrated substantial gains. Zhou et al. [17] formulated mmWave estimation as CP decomposition of a fully observed received signal tensor after RF combining—an overdetermined problem targeting hybrid architectures. Tensor methods have also found application in MIMO relay systems via nested Tucker decompositions [19] and in time-varying channels under high mobility [20]. However, these formulations assume fully observed tensors and do not address pilot-limited completion.
This paper addresses a complementary scenario: pilot-limited estimation in fully digital systems where only a sparse subset of channel tensor entries is observed. This severely underdetermined inverse problem requires regularization through low-rank tensor completion—a fundamentally different mathematical setting applicable to a broader class of MIMO-OFDM systems [4].
The main contributions of this article are:
@@
    Tensor completion formulation for pilot-limited MIMO: Unlike prior methods assuming fully observed tensors [17], this work formulates the problem as tensor completion from sparse pilot samples.
@@
    Hybrid tensor–neural architecture: A two-stage estimator combining low-rank tensor completion with neural residual learning, achieving 1–5 dB NMSE reduction over pure tensor-based methods on DeepMIMO ray-tracing channels, with the hybrid estimator remaining stable under extreme pilot scarcity where CP-based methods exhibit numerical instability.
@@
    Comprehensive validation: Evaluation against LS and OMP baselines demonstrates 7–11 dB NMSE improvement at 5–10% pilot density on synthetic channels, and up to 14 dB on DeepMIMO ray-tracing channels at comparable pilot densities.
@@ 2 System Model
This paper denotes scalars by lowercase letters (e.g., $\alpha$), vectors by bold lowercase letters (e.g., $\mathbf{a}$), matrices by bold uppercase letters (e.g., $\mathbf{A}$), and tensors by calligraphic letters (e.g., $\mathcal{H}$). The wideband MIMO channel takes the form of a third-order tensor $\mathcal{H}\in\mathbb{C}^{N_{r}\times N_{t}\times N_{f}}$, where $\mathbf{w}[k]$ denotes the $k$-th entry of a vector $\mathbf{w}$. The symbol $\circ$ denotes the outer product, $\times_{n}$ the $n$-mode product, $\|\cdot\|_{F}$ the Frobenius norm, $\odot$ the Hadamard (element-wise) product, $\otimes$ the Kronecker product, and $(\cdot)^{H}$ the conjugate transpose (adjoint). All tensor dimensions and ranks follow this convention throughout the paper.
Let $\mathbf{H}[k]\in\mathbb{C}^{N_{r}\times N_{t}}$ denote the frequency-domain channel matrix at subcarrier $k\in\{1,\ldots,N_{f}\}$. The received signal satisfies
@@ 2.1 Pilot-Based Channel Estimation
During the training phase, known pilot symbols occupy a subset of resource elements $(i,j,k)$ of the channel tensor, so observations cover only a subset of its coefficients. Stacking the frequency-domain channel matrices $\mathbf{H}[k]\in\mathbb{C}^{N_{r}\times N_{t}}$ across all $N_{f}$ subcarriers yields the wideband MIMO channel as a third-order tensor
$\mathcal{H}\in\mathbb{C}^{N_{r}\times N_{t}\times N_{f}}.$ (2)
Pilot-based acquisition covers only the subset of tensor entries indexed by
@@
The resulting observation model reads [17]
@@
and $\mathcal{N}$ is i.i.d. complex Gaussian noise. and $\mathcal{N}$ is i.i.d. complex Gaussian noise. In practical pilot-aided OFDM training, the channel tensor admits no direct observation. At each pilot location $(i,j,k)\in\Omega$, a known pilot symbol provides a noisy scalar observation of the corresponding channel coefficient $\mathcal{H}_{ijk}$. Accordingly, $\mathcal{P}_{\Omega}(\mathcal{H})$ collects these observations as entry-wise samples of the channel tensor corrupted by additive noise. This abstraction reduces the estimation problem to entry-wise sampling of the channel tensor with noise, which is standard in tensor completion and isolates the impact of pilot sparsity from waveform-specific effects. This article defines SNR with respect to the average power per channel coefficient over the full tensor $\mathcal{H}$, independent of the pilot set $\Omega$ and pilot ratio $\rho$, as formalized in (16). This ensures strict SNR comparability across all pilot densities. Equivalently, $\mathcal{P}_{\Omega}(\mathcal{H})=\mathcal{M}\odot\mathcal{H}$, where $\mathcal{M}\in\{0,1\}^{N_{r}\times N_{t}\times N_{f}}$ is a binary mask with $\mathcal{M}]_{ijk}=1$ if $(i,j,k)\in\Omega$. Stacking the per-subcarrier noise vectors $\mathbf{n}[k]$ across receive antennas and frequency indices yields the noise tensor $\mathcal{N}$, consistently with (1). This article defines SNR with respect to the average power per channel coefficient over the full tensor $\mathcal{H}$, independent of the pilot set $\Omega$ and pilot ratio $\rho$, as formalized in (16). This ensures strict SNR comparability across all pilot densities. Equivalently, $\mathcal{P}_{\Omega}(\mathcal{H})=\mathcal{M}\odot\mathcal{H}$, where $\mathcal{M}\in\{0,1\}^{N_{r}\times N_{t}\times N_{f}}$ is a binary mask with $\mathcal{M}]_{ijk}=1$ if $(i,j,k)\in\Omega$. Stacking the per-subcarrier noise vectors $\mathbf{n}[k]$ across receive antennas and frequency indices yields the noise tensor $\mathcal{N}$, consistently with (1).
The tensor completion formulation treats pilot-limited observation as entry-wise sampling over an index set $\Omega$. The experiments draw $\Omega$ from a pilot mask with either random sampling, a regular grid, or a comb pattern in frequency. Random uniform sampling serves as the default model, consistent with standard practice in tensor completion literature [26, 33]. Structured masks define deterministic index sets that satisfy the same observation model but may exhibit different empirical recovery behavior; their performance is assessed experimentally in Experiment 4.
@@
only a small fraction of channel tensor coefficients admits direct observation. In this pilot-limited regime, recovering the full channel tensor $\mathcal{H}$ from partial and noisy observations $\mathcal{Y}_{\Omega}$ constitutes a severely ill-posed inverse problem: unknowns vastly exceed measurements, the observation operator is highly rank-deficient, and the solution is non-unique and noise-sensitive [6, 7, 22]. This fundamental limitation motivates incorporating structural priors that exploit intrinsic spatial and spectral correlations of the channel to regularize the estimation problem and enable reliable reconstruction.
@@ 2.2 Multipath Channel Model
where $L$ denotes the number of dominant propagation paths, $\alpha_{\ell}\in\mathbb{C}$ are complex path gains, $\mathbf{a}_{r}(\theta_{\ell})$ and $\mathbf{a}_{t}(\phi_{\ell})$ are receive and transmit array steering vectors associated with angles of arrival and departure, respectively, assuming a 2D propagation model with azimuth-only angular parameters, and $\tau_{\ell}\in[0,1)$ represents the normalized propagation delay. The exponential term models frequency-dependent phase rotation induced by path delay across OFDM subcarriers.
Eq. (7) reveals that each propagation path contributes a separable outer-product structure across receive, transmit, and frequency modes. Consequently, the wideband channel tensor takes the form of a sum of rank-one components, yielding an intrinsic low-rank CP representation governed by the number of dominant paths [17, 23]. This physically grounded low-rank structure directly motivates tensor-based channel modeling and underpins the structure-informed estimation framework developed in Section 3.
@@ 3 Tensor Representations and Algebraic Priors
This section formalizes how fundamental propagation mechanisms in wideband MIMO channels manifest as intrinsic algebraic structure. It introduces CP and Tucker decompositions as principled parameterizations of this structure, providing physically interpretable and computationally efficient priors for pilot-limited channel estimation.
@@ 3.1 Canonical Polyadic Decomposition
where $k$ indexes the subcarrier frequencies. Absorbing the complex gain $\alpha_{\ell}$ into the receive-mode factor $\mathbf{u}_{\ell}$ is a matter of convention; the gain could equivalently enter $\mathbf{v}_{\ell}$, $\mathbf{w}_{\ell}$, or distribute across multiple factors. This scaling ambiguity is intrinsic to the CP decomposition: under the Kruskal uniqueness condition, the CP model is unique only up to permutation and scaling of its rank-one components [23].
This correspondence makes explicit that each rank-one tensor component represents a single physical propagation path, implying that the effective CP rank is governed by the number of dominant paths rather than the ambient channel dimensionality. As a result, CP is particularly well matched to propagation regimes dominated by specular ray-like components, such as millimeter-wave channels around 28 GHz, where a small number of strong paths typically captures most of the channel energy. This property underlies the effectiveness of CP-based models for sparse-channel and large-scale MIMO estimation [17]. The Kruskal uniqueness condition guarantees factor identifiability up to permutation and scaling, which suffices for reliable channel reconstruction. Extracting explicit physical parameters, such as angles of arrival or departure, from the estimated factors would require additional structural constraints—e.g., array calibration or explicit enforcement of Vandermonde structure—beyond the scope of this work, which focuses on CSI reconstruction rather than direct parameter estimation [16].